# thanks: Talento-CM fellow

### v1

@@ III.2 $\nu w_{0}w_{a}$ model
Figure 3: Similar to Fig. 2, but in $\nu w_{0}w_{a}$ model.
Next we extend the $\nu\Lambda$CDM model to the $\nu w_{0}w_{a}$ model in order to investigate the impact of the dynamic dark energy on the neutrino constraint, referring to columns 8-13 in Table 2. The corresponding contour plot is shown in Fig. 3 for the cosmological parameters $\{M_{\nu},w_{0},w_{a}\}$ with other parameters marginalized over.
It is expected that including the parameters $\{w_{0},w_{a}\}$ will loosen the constraint on $M_{\nu}$ since both dark energy and massive neutrinos influence the growth rate of structures in the Universe. As depicted in Fig. 3, there is a small positive correlation between $M_{\nu}$ and $w_{0}$, as well as a distinct negative correlation between $M_{\nu}$ and $w_{a}$, for both $N$ and $P$. However, the direction of degeneracy is slightly different between $N$ and $P$, resulting in the constraint from the combination $N+P$ being less affected by the presence of the dynamic dark energy [73].
Consider the known SR scenario as an example. Adding $\{w_{0},w_{a}\}$ results in a loosening of $\sigma(M_{\nu})$ by a factor of 1.552 and 1.1471 for $N$ and $P$, respectively. In contrast, the combination $N+P$ demonstrates a loosening factor of only 1.079. The same holds true for the self-calibration scenario, where the loosening factor for the combination $N+P$ is 1.084.
In summary, the constraint on the total neutrino mass, $M_{\nu}$, is mainly contributed by the cluster power spectrum, $P$. The inclusion of dynamic dark energy has minimal impact, whereas the availability of perfect SR knowledge plays a significant role.
@@
Table 3: Forecasted the $1\,\sigma$ marginalized error for the the total neutrino for different conditions. The 4th row represents the fiducial result, same as shown in Table 2. The subsequent rows demonstrate the results under specific conditions: partial SR knowledge (5th row), $z_{max}=2$ (6th row), $\sigma_{z}=0.01$ (7th row), HMF uncertainty (8th row), bias uncertainty (9th row), and without considering the GISDB effect (10th row). The symbol ’/’ denotes that the constraint is the same as the fiducial condition.
@@
In this section, we analyze several systematic effects and survey strategies, as presented in Table 3, and conclude by comparing our forecast with results from other surveys at the end of the section.
@@ IV.1 SR parameters
In our analysis, we do not directly incorporate the mass observable in CSST, i.e. the optical richness, into our model. Although the mass-richness relation (MRR) has been calibrated in numerous works (e.g. [32, 33, 34, 35, 36]), it will be affected by different surveys and cluster finders [74, 75]. Therefore it is inappropriate to select a MRR calibrated in one specific survey to another survey. Instead, we employ a simple parameterization to account for our current lack of knowledge regarding the details of the mass-observable relations and their potential dependence on redshift and mass. However, our approach is not entirely devoid of information. Rather than treating $\sigma_{\ln M}$ as a nuisance parameter, we relate it with $\sigma_{\ln\lambda}$, and assume $\sigma_{\ln\lambda}$ comprises a combination of a constant intrinsic scatter and a Poisson-like term. In our main analysis, we have selected a MRR to model the Poisson-like term $1/\exp\left<\ln\lambda^{ob}(M)\right>$. However, in actual surveys, we will utilize a specific MRR to replace Eq.(10), and there will be no need for an additional MRR to account for the Poisson-like term $1/\exp\left<\ln\lambda^{ob}(M)\right>$. Consequently, we fix parameters $\{A_{\lambda},B_{\lambda},C_{\lambda}\}$ in Eq.(II.1), while retaining $B_{\lambda}$ as a nuisance parameter in Eq.(14) to mimic this scenario, which we refer to as the partial self-calibration scenario.
The 5th row in Table 3 shows the result. Similar to the previous section, SR parameters have a greater impact on $N$ than on $P$. The partial self-calibration scenario improve the constraint on $M_{\nu}$ from $N$ by a factor of 1.8 (1.2) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model. However, the improvement is minimal for the constraint from $P$ or $N+P$.
@@
Recently, we calibrate the intrinsic mass-richness relation of clusters in THE THREE HUNDRED hydrodynamic simulations [76]. Our findings indicate that employing a skewed Gaussian distribution to model the richness at a fixed halo mass results in a simpler and smaller scatter compared to the conventional use of a log-normal distribution. Once this scaling relation is validated with the CSST mock data, we intend to incorporate it into future cosmological analyses.
@@ IV.2 $z_{max}$
In our fiducial analysis, we assume a maximum redshift of $z_{\max}=1.5$ for cluster detection with CSST. However, by improving the cluster selection algorithm or jointing analysis with other surveys like Euclid [31], it may be possible to extend the maximum redshift to $z_{\max}=2$ [30]. The 6th row in Table 3 shows the result with $z_{max}=2$. Compared to the fiducial result, the total number counts is increased to 434534, and the constraint from $N$ has been improved by a factor of $\sim$1.7 (1.1) in the self-calibration (known SR) scenario for both two cosmological models. However, the improvement for the constraint from $P$ is very limited, with a factor of only $\sim$1.02 across all cases. Since the constraint on $M_{\nu}$ primarily comes from $P$, the improvement in $\sigma^{N+P}(M_{\nu})$ is not significant. In the self-calibration (known SR) scenario, the improvement is $\sim$1.1 (1.03) for both cosmological models.
We also investigate a wide range of $z_{max}$, as depicted in Fig. 5. The upper panel shows the $\sigma(M_{\nu})$ in $\nu\Lambda$CDM model. It is evident that $\sigma^{N}(M_{\nu})$ decreases significantly with the increase of $z_{max}$. Conversely, $\sigma^{P}(M_{\nu})$ remains relatively constant once $z_{max}$ reaches $\sim$1.4. This constancy arises due to the exponential decrease in the cluster number density $\bar{n}$ Eq.(31) with increasing redshift, resulting in a very small effective volume $V_{\text{eff}}$ Eq.(II.4) at high redshifts. Consequently, the constraint from the combination $N+P$, i.e. $\sigma^{N+P}(M_{\nu})$, exhibits a moderate decrease with increasing $z_{\text{max}}$.
@@
Figure 6: The impact of the redshift uncertainty $\sigma_{z}$ on $\sigma(M_{\nu})$ in $\nu w_{0}w_{a}$ CDM model. Different colors (red, blue and black) represent different probes: the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$. Solid and dotted lines represent self-calibration and known SR scenario respectively.
The redshift uncertainty $\sigma_{z}$ is optimistically assumed to be 0.001 in our fiducial analysis. This assumption is based on the availability of spectroscopic redshifts for all clusters observed by the CSST. However, in cases where this assumption does not hold, we consider a more conservative estimate of 0.01 for $\sigma_{z}$. The constraint result using this conservative estimate is presented in the 7th row of Table 3. We also extend $\sigma_{z}$ up to 0.03 in Fig. 5, which is the photometric redshift error for Euclid at high redshift [72].
The redshift uncertainty $\sigma_{z}$ has a minimal impact on $\sigma^{N}(M_{\nu})$ due to the redshift bin size. Even for the largest redshift uncertainty $\sim$0.03, it is still smaller than the redshift bin $\Delta z=0.05$ we choose. However, the influence of $\sigma_{z}$ on $\sigma^{P}(M_{\nu})$ is noteworthy. This is attributed to the damping effect caused by the redshift uncertainty, which leads to the partial erasure of information in $P$. It is observed that the behavior of $\sigma^{N+P}(M_{\nu})$ aligns with that of $\sigma^{P}(M_{\nu})$. Quantitatively, the ratio between $\sigma^{N+P}(M_{\nu})$ with $\sigma_{z}=0.01$ and $\sigma_{z}=0.001$ is $\sim$2. These findings underscore the significance of precise redshift measurements in effectively constraining the total neutrino mass.
@@ IV.4 HMF uncertainty
Although the CDM prescription we use in subsection II.3 has alleviate the impact of neutrinos on the universality of HMF, the HMF of Tinker we employ is accurate at the $\sim$5level [26]. Additionally, baryonic effects introduce an uncertainty of $\sim$10to the HMF [78, 79]. To account for these uncertainties, we incorporate two additional parameters $\{s,q\}$ to characterize the deviation of the Tinker’s HMF from the true HMF, following [80] and [81]:
@@
The fiducial values are $s=0$ and $q=1$, degenerating the HMF to the Tinker’s.
The 8th row in Table 3 presents the constraints with additional HMF uncertainties. Taking the $\nu w_{0}w_{a}$ model as an example, compared to $\sigma(M_{\nu})$ in the fiducial condition, the constraints from $N$ have been relaxed by a factor of 1.2 (2.7) in the self-calibration (known SR) scenario, while the constraints from $P$ remain largely unchanged. However for the constraints from $N+P$, the relaxation becomes more significant by a factor of 1.85 (3.97). Similar conclusions hold for the $\nu\Lambda$CDM model. As a result, the accuracy of the HMF plays a crucial role in constraining $M_{\nu}$ using $N+P$. It is worth mentioning that current and future emulators, based on simulations with massive neutrinos, have the potential to determine the HMF with greater precision, bypassing the reliance on a universal formula [82].
IV.5 bias uncertainty
Galaxy clusters are biased tracers of the CDM field. Fortunately, compared to galaxies, we have theory to describe the bias $b(M,z)$ in Eq.(17) for clusters, and the formula can be calibrated through simulations, similar to the HMF. Therefore, the bias of Tinker that we utilize also exhibits uncertainties. Moreover, in real surveys, extra bias will be introduced by selection effects, primarily due to projection effects and orientation biases. We incorporate two parameters $\{b_{s0},b_{s1}\}$ to model these uncertainties, following [83, 84]:
@@
The constraint on $M_{\nu}$ with the bias uncertainty is presented in the 9 th row in the Table 3. For $\sigma^{P}(M_{\nu})$, in the self-calibration scenario, the effect of bias uncertainty is overshadowed by the effect of SR parameters. However, in the known SR scenario, the inclusion of bias uncertainty relaxes the constraint by a factor of $\sim$1.03 (1.3) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model. On the other hand, for $\sigma^{N+P}(M_{\nu})$, even in the self-calibration scenario, the effect of bias uncertainty is significant with a factor of $\sim$1.38(1.96). In the known SR scenario, it is 2.94 (3.5) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model. These findings highlight the greater influence of bias uncertainty compared to that of the HMF. It emphasizes the need for careful consideration of bias uncertainties in future analyses.
@@ IV.6 GISDB
Figure 7: Bias. The impact of the GISDB effect on $\sigma(M_{\nu})$ in $\nu w_{0}w_{a}$ CDM model. Different colors (blue and black) represent different probes: cluster power spectrum $P$ and the combination $N+P$. Solid and hollow dots represent results with or without GISDB effect. Each column represents the self-calibration scenario (SC), the known SR scenario (SR), SC with the bias uncertainty and SR with the bias uncertainty, respectively.
In this section, we investigate the impact of the GISDB effect, which has been omitted in previous studies (e.g. [72, 31, 85]). The constraint without GISDB is presented in the 10th row of Table 3.
GISDB effect compensates for the suppression of the power spectrum caused by neutrinos at small scales.As a result, we anticipate that the constraint obtained from $P$ will be relaxed when GISDB is considered. However, there is an exception, $\sigma^{P}(M_{\nu})$ changing from 0.229 to 0.223. The reason for this exception is that when using $P$ to constrain $M_{\nu}$, we utilize information not only from the shape of the power spectrum but also from its amplitude, predominantly through the effective bias $b_{\text{eff}}$. By adding bias uncertainties, as discussed in the previous subsection, the amplitude information is blurred, leaving only the shape information.
@@
IV.7 Comparison with other cluster surveys
@@
In this subsection, we compare our results for CSST with others for LSST, Euclid, and CMB-S4. All the constraints discussed here are derived in the $\nu w_{0}w_{a}$ model, except for LSST, which is in the $\nu w_{0}$ model.
@@
The first work using cluster number counts and cluster power spectrum to constrain the total neutrino mass was performed by Wang et al. for LSST, a shear-limited weak lensing survey with a sky area of 18,000 deg2 and a redshift range of $z=[0.1,1.4]$. They forecast $\sigma^{P}(M_{\nu})=0.71$ eV and $\sigma^{N+P}(M_{\nu})=0.42$ eV, with an improvement factor of $\sim$1.34 when adding $N$. In the self-calibration scenario, we obtain $\sigma^{P}(M_{\nu})=0.293$ eV and $\sigma^{N+P}(M_{\nu})=0.127$ eV, with a improvement factor of $\sim$2.31. The main difference between our results and those of Wang et al. lies in the improvement factors of 1.34 and 2.31, respectively. This disparity can be attributed to the fact that the total cluster number counts forecasted by Wang et al. is approximately $2\times 10^{5}$, which is lower than our estimate of $4\times 10^{5}$.
@@
Next, we compare our results with those obtained by Sartoris et al. for Euclid, an optical survey with a sky area of 15,000 deg2 and a redshift range of $z=[0.2,2]$ [31]. They use a selection threshold of $N_{500,c}/\sigma_{\text{field}}$, a ratio of $N_{500,c}$ the number of cluster galaxies contained within a sphere of certain radius and $\sigma_{\text{field}}$ the rms of the field galaxy counts. They forecast a total cluster number counts of $2\times 10^{6}$ and $2\times 10^{5}$ for $N_{500,c}/\sigma_{\text{field}}=3$ and 5 respectively. Our $4\times 10^{5}$ lies between them. For the former selection $N_{500,c}/\sigma_{\text{field}}=3$, they forecast $\sigma^{N+P}(M_{\nu})=0.140$ eV in the self-calibration scenario. When in the known SR scenario, it becomes 0.121 with an improvement factor of $\sim$1.16. Our main result is $\sigma^{N+P}(M_{\nu})=0.127$ eV in self-calibration, slightly smaller than $0.140$ eV. However, perfect knowledge of SR parameters bring our constraint greater improvement of $\sim$3.7 to $\sigma^{N+P}(M_{\nu})=0.034$ eV. When extending $z_{max}$ from 1.5 to 2, our results become smaller $\sigma^{N+P}(M_{\nu})=0.118\ (0.033)$ eV in self-calibration (known SR) scenario. It is worth noting that if we do not consider the GISDB effect as they do, our result becomes $\sigma^{N+P}(M_{\nu})=0.191\ (0.075)$ eV, larger than their $0.140$ eV.
@@
Finally, we compare our results with those obtained by Cromer et al. for CMB-S4, a Sunyaev-Zel’dovich (SZ) survey with a sky area of 10,000 deg2 and a redshift range of $z=[0.1,1.9]$ [85]. They forecast a total cluster number counts of $\sim 10^{5}$. By combining Planck primary CMB and extra mass calibration from CMB lensing, they forecast $\sigma^{N+P}(M_{\nu})\sim 0.1$ eV, which falls between our results of $0.127$ eV and $0.034$ eV in the self-calibration and known SR scenarios, respectively. It is worth noting that they find the improvement on $\sigma(M_{\nu})$ when adding $P$ is very small $\sim$0.5%. This may be due to the fact that they have already incorporated prior information from the primary CMB and additional mass calibration, resulting in the $N$ alone having a certain constraining ability. Furthermore, their redshift uncertainty is reported to be 0.01, which has a significant impact on the constraining ability of $P$. They also consider the bias uncertainty and find it has less than 1% effect on $\sigma(M_{\nu})$, mainly because $P$ has a minimal contribution on the constraint.
@@
In this work, we forecast constraints on the total neutrino mass from CSST galaxy clusters. Specifically, we employ Fisher matrix analysis to derive the constraint $\sigma(M_{\nu})$ from the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$. In addition to considering the most basic cosmological extension, the standard cosmology plus neutrinos, $\nu\Lambda$CDM model, we also explore the inclusion of dynamic dark energy, $\nu w_{0}w_{a}$ model. Since the largest source of systematics in galaxy cluster cosmology arises from the mass-observable relation, we examine cases where the SR parameters are completely unknown and perfectly known as self-calibration scenario and known SR scenario. The main results are presented in Table 3. The results between $\nu\Lambda CDM$ and $\nu w_{0}w_{a}$ model are similar. Unless otherwise specified, the following results refer to $\nu w_{0}w_{a}$ model:
@@
  • CSST is expected to detect 408037 clusters with the mass threshold $M_{200m}\geq 0.836\times 10^{14}M_{\odot}/h$ and a survey area of 17,500 deg2 in the redshift range of $z=[0,1.5]$ , as depicted in Fig. 1.

### v2

@@ III.2 $\nu w_{0}w_{a}$ model
Figure 3: Similar to Fig. 2, but for the $\nu w_{0}w_{a}$ model.
Next we extend the $\nu\Lambda$CDM model to the $\nu w_{0}w_{a}$ model in order to investigate the impact of dynamic dark energy on the neutrino constraint, results showing in columns 8-13 in Table 2. The corresponding contour plot is shown in Fig. 3 for the cosmological parameters $\{M_{\nu},w_{0},w_{a}\}$ with the other parameters marginalized over.
It is expected that including the parameters $\{w_{0},w_{a}\}$ will weaken the constraint on $M_{\nu}$ since both dark energy and massive neutrinos affect the growth rate of structures in the Universe. As depicted in Fig. 3, there is a positive correlation between $M_{\nu}$ and $w_{0}$, as well as a negative correlation between $M_{\nu}$ and $w_{a}$, for both $N$ and $P$. However, the direction of degeneracy is different between $N$ and $P$, resulting in the constraint on $M_{\nu}$ from the combination $N+P$ being less affected by the presence of the dynamic dark energy [77].
Consider the known SR scenario as an example. Adding $\{w_{0},w_{a}\}$ results in a weakening of $\sigma(M_{\nu})$ by a factor of 1.55 and 1.15 for $N$ and $P$, respectively. In contrast, the combination $N+P$ shows a weakening factor of only 1.10. The same holds true for the self-calibration scenario, where the weakening factor for the combination $N+P$ is 1.08.
In summary, the total neutrino mass, $M_{\nu}$, is much better constrained by the cluster power spectrum than by the cluster number counts. The two probes are highly complementary in constraining $M_{\nu}$. The inclusion of dynamic dark energy has negligible impact on the joint neutrino mass constraint from $N+P$. Availability of perfect SR knowledge plays a significant role in constraining $M_{\nu}$.
@@
Table 3: The $1\,\sigma$ marginalized errors forecasted for the the total neutrino mass under different conditions. The 4th row give the fiducial results, as shown in Table 2. The following rows show the results under the conditions of: partial SR knowledge (5th row), $z_{max}=2$ (6th row), $\sigma_{z}=0.01$ (7th row), HMF uncertainty (8th row), bias uncertainty (9th row), and without considering the GISDB effect (10th row). The symbol ‘/’ represents that the constraint is the same as the fiducial case.
@@
In this section, we study several effects that may change our forecasts, including knowledge on SR parameters, the maximum redshift, uncertainties in redshift measurements, HMF and bias, as well as the GISDB effect, with results presenting in Table 3, and compare with literature at the end of this section.
@@ IV.1 SR parameters
In our analysis, we do not directly use the mass observable for CSST, i.e. the optical richness. Although the mass-richness relation (MRR) has been calibrated in numerous works (e.g. [32, 33, 34, 35, 36]), it will be affected by survey specifics and cluster finders [78, 79]. Therefore it is inappropriate to apply a MRR calibrated in one specific survey to another survey. Instead, we employ a simple parametrization to account for our current lack of knowledge about the details of the mass-observable relations and their potential dependences on redshift and mass. However, our approach is not completely devoid of information. Rather than treating $\sigma_{\ln M}$ as a nuisance parameter, we relate it with $\sigma_{\ln\lambda}$, and assume $\sigma_{\ln\lambda}^{2}$ to be composed of a constant intrinsic scatter term and a Poisson-like term. In our default analysis, we have selected a MRR to model the Poisson-like term of $1/\exp\left<\ln\lambda^{ob}(M)\right>$. However, in real surveys, we can also use the richnesses for the detected clusters directly, then there will be no need for an additional MRR to specify the Poisson-like term. Consequently, we try fixing the parameters $\{A_{\lambda},B_{\lambda},C_{\lambda}\}$ in Eq.(II.1) while retaining $B_{\lambda}$ as a free parameter in Eq.(14) to describe this scenario, which we refer to as the partial self-calibration scenario.
The 5th row in Table 3 shows the result. Similar to the previous section, SR parameters have a greater impact on $N$ than on $P$. In the partial self-calibration scenario, the constraint on $M_{\nu}$ is improved by a factor of 1.8 (1.2) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model for $N$. However, the improvement is minimal for the constraint from $P$ or $N+P$.
@@
Recently, we have calibrated the intrinsic mass-richness relation of clusters with THE THREE HUNDRED hydrodynamic simulations [80]. Our findings indicate that employing a skewed Gaussian distribution to model the richness at a fixed halo mass results in a more accurate compared to the conventional use of a log-normal distribution. We plan to incorporate these findings in future cluster analyses.
@@ IV.2 $z_{max}$
In our fiducial analysis, we assume a maximum redshift of $z_{\max}=1.5$ for cluster detection with CSST. However, by improving the cluster selection algorithm or combining data from other surveys like Euclid [31], it may be possible to extend $z_{\max}=1.5$ to be even higher such as to $z_{\max}=2$ [30]. The 6th row in Table 3 shows the result with $z_{max}=2$. Compared to the fiducial result, the total number count is increased to 434534, and the constraint from $N$ has been improved by a factor of $\sim$1.7 (1.1) in the self-calibration (known SR) scenario for both cosmological models. However, the improvement for the constraint from $P$ is very limited, with a factor of only $\sim$1.02 in all cases, probably due to the large shot noises in the higher redshift bins. Since the constraint on $M_{\nu}$ is dominated by $P$, the improvement in $\sigma^{N+P}(M_{\nu})$ is not significant either. In the self-calibration (known SR) scenario, the improvement is only $\sim$1.1 (1.03) for both cosmological models.
We also investigate a wide range of $z_{max}$, results showing in Fig. 5. The upper panel shows $\sigma(M_{\nu})$ in $\nu\Lambda$CDM model. $\sigma^{N}(M_{\nu})$ in the self-calibration scenario is too large to be shown in this figure, but it decreases significantly with the increase of $z_{max}$. Conversely, $\sigma^{P}(M_{\nu})$ remains relatively a constant after $z_{max}$ has reached $\sim$1.4. This constancy arises from the sharp drop in the cluster number density $\bar{n}$ Eq.(31) with increasing redshift, resulting in a very small effective volume $V_{\text{eff}}$ Eq.(II.4) at high redshifts. Consequently, the constraint from the combination $N+P$, i.e. $\sigma^{N+P}(M_{\nu})$, exhibits a much more moderate decrease with increasing $z_{\text{max}}$.
@@ IV.3 Redshift uncertainty
Figure 6: The impact of redshift uncertainty parameter $\sigma_{z}$ on $\sigma(M_{\nu})$ for $\nu w_{0}w_{a}$ CDM model. Different colors represent different probes: red for the cluster number counts $N$, blue for cluster power spectrum $P$, and black for their combination $N+P$. Solid and dotted lines represent self-calibration and known SR scenarios respectively.
The redshift uncertainty parameter $\sigma_{z}$ is optimistically assumed to be 0.001 in our fiducial analysis. This accuracy probably highly rely on the availability of accurate spectroscopic redshifts for all clusters observed by the CSST. In case it will be challenging to achieve such an accuracy, we consider a more conservative estimate of 0.01 for $\sigma_{z}$. The constraint using this conservative estimate is presented in the 7th row of Table 3. We also extend $\sigma_{z}$ up to 0.03 with results displayed in Fig. 6 for the $\nu w_{0}w_{a}$ specifically.
Similarly to Fig. 5, $\sigma^{N}(M_{\nu})$ in the self-calibration scenario is too large to be displayed in this figure. For both scenarios, the redshift uncertainty $\sigma_{z}$ has a minimal impact on $\sigma^{N}(M_{\nu})$ due to the relatively large redshift bin size. However, the influence of $\sigma_{z}$ on $\sigma^{P}(M_{\nu})$ is noteworthy. This is attributed to the damping effect caused by the redshift uncertainty, which suppresses $P$ on relatively smaller scales. The behavior of $\sigma^{N+P}(M_{\nu})$ with increasing $\sigma_{z}$ is similar to that of $\sigma^{P}(M_{\nu})$, but less significant. Quantitatively, the ratio between $\sigma^{N+P}(M_{\nu})$ with $\sigma_{z}=0.01$ and $\sigma_{z}=0.001$ is $\sim$2. These findings underscore the importance of precise redshift measurements in effectively constraining the total neutrino mass.
@@ IV.4 HMF uncertainty
Although the CDM prescription we adopt in subsection II.3 has alleviated the impact of massive neutrinos on the universality of HMF, the HMF of Tinker et al (2008) we employ is accurate at the $\sim$5level [26]. Additionally, systematics such as different halo finders, non-$\Lambda$CDM simulations, baryonic effects etc. introduce extra uncertainty to the HMF [44, 82, 83]. To account for these uncertainties, we incorporate two additional parameters $\{s,q\}$ to characterize the deviation of the Tinker’s HMF from the true HMF, following [84] and [85]:
@@
The fiducial values are $s=0$ and $q=1$, reducing to the Tinker’s mass function.
The 8th row in Table 3 shows the constraints with additional HMF uncertainties. Taking the $\nu w_{0}w_{a}$ model as an example, compared to $\sigma(M_{\nu})$ in the fiducial case, the constraints from $N$ have been relaxed by a factor of 1.2 (2.7) in the self-calibration (known SR) scenario, while the constraints from $P$ remain largely unchanged. However for the constraints from $N+P$, the degradation becomes more significant by a factor of 1.85 (3.97). For the $\nu\Lambda$CDM model, $\sigma(M_{\nu})$ from $N+P$ degrades by a factor of 1.09(2.13). As a result, the accuracy of the HMF plays a crucial role in constraining $M_{\nu}$ using $N+P$. It is worth mentioning that emulators for HMF, based on simulations with massive neutrinos, have the potential to achieve greater precision, bypassing the reliance on a universal formula [86].
IV.5 Bias systematics
Galaxy clusters are biased tracers of matter field. Fortunately, compared to galaxies, we have theory to describe the bias for clusters, and the formula can be calibrated with simulations, similar to the HMF. The bias of Tinker et al (2010) that we utilize already exhibits uncertainties. Moreover, in real surveys, extra bias will be introduced by cluster selection effects. For example, the redMaPPer clusters show an additional mass-dependent bias due to projection effects and orientation biases[87]. We introduce two parameters $\{b_{s0},b_{s1}\}$ to model systematics on cluster bias, following [87, 88]:
@@ IV.5 Bias systematics
The constraint on $M_{\nu}$ with the bias uncertainty is presented on the 9 th row of Table 3. For $\sigma^{P}(M_{\nu})$, in the self-calibration scenario, the effect of bias uncertainty is completely negligible. However, in the known SR scenario, the inclusion of bias uncertainty relaxes the constraint by a factor of $\sim$1.03 (1.3) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model. On the other hand, for $\sigma^{N+P}(M_{\nu})$, even in the self-calibration scenario, the effect of bias uncertainty is significant with a factor of $\sim$1.38(1.96). In the known SR scenario, it is 2.94 (3.5) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model. These findings highlight the importance of bias systematics to cluster cosmology, and emphasize the need for careful study of bias systematics in future analyses.
@@ IV.6 GISDB
Figure 7: The impact of the GISDB effect on $\sigma(M_{\nu})$ in $\nu w_{0}w_{a}$ CDM model. Different colors (blue and black) represent different probes: cluster power spectrum $P$ and the combination $N+P$. Solid and hollow dots represent results with or without GISDB effect. From left to right, the columns represent the self-calibration scenario (SC), the known SR scenario (SR), SC with the bias uncertainty and SR with the bias uncertainty, respectively.
In this subsection, we investigate the impact of the GISDB, which has been missed in previous cluster studies (e.g. [76, 31, 89]). The constraint without GISDB is presented in the 10th row of Table 3.
GISDB compensates partly for the suppression of the power spectrum caused by massive neutrinos on small scales. As a result, we expect that the constraint obtained from $P$ will be degraded when GISDB is considered. This is indeed what we find except for the “known SR” case in the $\nu w_{0}w_{a}$ model, where $\sigma^{P}(M_{\nu})$ gets slightly better from 0.229 to 0.223. When using $P$ to constrain $M_{\nu}$, we utilize information not only from the shape of the power spectrum but also from its amplitude, predominantly through the effective bias $b_{\text{eff}}$. By adding bias uncertainties, as discussed in the previous subsection, the amplitude information is blurred, leaving only the shape information.
@@
IV.7 Comparison with other forecasts
@@
In this subsection, we compare our results forecasted for the CSST with other works in the literature forecasted for e.g. the LSST, Euclid, and CMB-S4. All the constraints discussed here are derived for the $\nu w_{0}w_{a}$ model, except for LSST, which is for the $\nu w_{0}$ model.
@@
The first work using cluster number counts and cluster power spectrum to constrain the total neutrino mass was performed by Wang et al. (2005) for several cluster surveys including the LSST, which can provide a shear-selected cluster catalogue in a sky area of 18,000 deg2 and a redshift range of $z=[0.1,1.4]$. They forecast $\sigma^{P}(M_{\nu})=0.71$ eV and $\sigma^{N+P}(M_{\nu})=0.42$ eV, with an improvement factor of $\sim$1.70 when adding $N$. In the self-calibration scenario, we obtain $\sigma^{P}(M_{\nu})=0.293$ eV and $\sigma^{N+P}(M_{\nu})=0.127$ eV, with a improvement factor of $\sim$2.31. This difference can be attributed to the fact that the total cluster number counts forecasted by Wang et al. is approximately $2\times 10^{5}$, which is much lower than our estimate of $4\times 10^{5}$.
@@
Next, we compare our results with those obtained by Sartoris et al (2016) forecasted for the Euclid, which can provide an optical cluster catalogue with a sky area of 15,000 deg2 and a redshift range of $z=[0.2,2]$ [31]. They use a selection threshold of $N_{500,c}/\sigma_{\text{field}}$, a ratio of $N_{500,c}$ the number of cluster galaxies contained within a certain radius to $\sigma_{\text{field}}$ the rms of the field galaxy counts. They forecast a total cluster number counts of $2\times 10^{6}$ and $2\times 10^{5}$ for $N_{500,c}/\sigma_{\text{field}}=3$ and 5 respectively. Our cluster counts $4\times 10^{5}$ lies between these two numbers. For the former selection $N_{500,c}/\sigma_{\text{field}}=3$, they forecast $\sigma^{N+P}(M_{\nu})=0.140$ eV in the self-calibration scenario. In the known SR scenario, it becomes 0.121eV with an improvement factor of $\sim$1.16. Our main result is $\sigma^{N+P}(M_{\nu})=0.127$ eV in self-calibration, slightly smaller than $0.140$ eV. However, perfect knowledge of SR parameters brings our constraint greater improvement by $\sim$3.7 to $\sigma^{N+P}(M_{\nu})=0.034$ eV. When extending $z_{max}$ from 1.5 to 2, our results become even smaller $\sigma^{N+P}(M_{\nu})=0.118\ (0.033)$ eV in self-calibration (known SR) scenario. It is worth noting that if we do not consider the GISDB effect as they do, our result becomes $\sigma^{N+P}(M_{\nu})=0.191\ (0.075)$ eV, larger than their $0.140$ eV.
@@
Finally, we compare our results with forecasts by Cromer et al (2019) for the CMB-S4, which can provide a Sunyaev-Zel’dovich (SZ) cluster catalogue with a sky area of 10,000 deg2 and a redshift range of $z=[0.1,1.9]$ [89]. They forecast a total cluster number counts of $\sim 10^{5}$. By combining with Planck primary CMB anisotropies and extra mass calibration from weak lensing, they obtain $\sigma^{N+P}(M_{\nu})\sim 0.1$ eV, which falls between our results of $0.127$ eV and $0.034$ eV in the self-calibration and known SR scenarios, respectively. It is worth noting that they find the improvement on $\sigma(M_{\nu})$ when adding $P$ is very small $\sim$0.5%. This may be due to the fact that they have already incorporated information from the primary CMB and additional mass calibration in cluster number counts, making adding in $P$ less effective. Furthermore, their redshift uncertainty is set to be a conservative value of 0.01, which has a significant impact on the constraining ability of $P$. They also consider the bias uncertainty and find it has negligible effect on the combined constraint on $\sigma(M_{\nu})$, mainly because $P$ has minimal contribution to the constraint.
@@
In this work, we forecast constraints on the total neutrino mass from the CSST galaxy clusters. Specifically, we employ the Fisher matrix formalism to derive the constraint from the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$. In addition to considering the simplest extension to the $\Lambda$CDM model, the $\nu\Lambda$CDM model ($\Lambda$CDM plus massive neutrinos), we also explore the inclusion of dynamic dark energy, the $\nu w_{0}w_{a}$ model. We carefully take into account the effects of massive neutrinos on the halo mass function and halo bias. Since the largest source of systematics in galaxy cluster cosmology arises from the mass-observable relation, we examine two cases where the SR parameters are completely unknown or perfectly known, named as the self-calibration scenario and known SR scenario respectively. Our main results are presented in Table 2 and Table 3. The results for the $\nu\Lambda CDM$ and $\nu w_{0}w_{a}$ model are similar. Unless otherwise specified, the following results refer to the $\nu w_{0}w_{a}$ model:
@@
  • CSST is expected to detect $\sim 400,000$ clusters with a mass threshold of $M_{200m}=0.836\times 10^{14}M_{\odot}/h$ in a survey area of 17,500 deg2 and redshift range of $z=[0,1.5]$ , as depicted in Fig. 1.